OpenStudy (anonymous):

Calculate the pH of the following solution: A sample of seawater that has an OH^- concentration of 1.58x10^-6 M

OpenStudy (imtiaz7):

pH+pOH =14 pH=14-pOH pH=14-1.58*10^-6 pH=12.42*10^-6